# Advancements in Constitutive Model Calibration: Leveraging the Power of Full-Field DIC Measurements and In-Situ Load Path Selection for Reliable Parameter Inference

### v1

@@ 3.4 Bayesian Techniques
A brief discussion of the Bayesian paradigm for inference and experimental design is contained in this section. The reader is referred to [27] for a complete discussion on the theory and tools found in this work.
@@ 3.4.1 Bayesian Inference
In each cycle of the ICC feedback loop, a single load step is applied to the specimen, and the collected data is used to perform model calibration. In lieu of a deterministic approach for model calibration, Bayesian inference is used, which provides an elegant way to quantify parameter uncertainty.
The first step is to model any knowledge regarding the unknown parameters of the forward model, $\boldsymbol{\theta}\in\boldsymbol{\Theta}\subset\mathbb{R}^{D}$, that is available a priori to observing data. In the adopted notation, the parameters $\boldsymbol{\theta}$ belongs to subspace $\boldsymbol{\Theta}$ of dimension $D$. Data $\boldsymbol{y}\in\boldsymbol{Y}\subset\mathbb{R}^{N_{obs}}$ is then collected which belongs to the data space $\boldsymbol{Y}$ with dimensionality $N_{obs}$, which is the number of measurements. The probability model for the data is denoted by $f\left(\boldsymbol{y}\>|\>\boldsymbol{\theta}\right)$, which describes the probability of observing data $\boldsymbol{y}$ given parameters $\boldsymbol{\theta}$, and for fixed $\boldsymbol{y}$, is regarded as the likelihood function. Once this data become available, the prior model is conditioned upon it and updated to the posterior distribution, $\pi(\boldsymbol{\theta}\>|\>\boldsymbol{y})$, via Bayes’ rule,
@@
Thus, both the data and prior beliefs are taken into account when calculating the posterior distribution, which is, generally speaking, not available in closed-form due to the dimensionality of the parameter space. For this reason, simulation techniques such as Markov Chain Monte Carlo (MCMC) [58, 59, 60, 61, 62] are widely used to obtain an approximation of the posterior distribution through one or more of a variety of sampling algorithms. However, in practical applications for real-time decision-making, MCMC is prohibitively expensive. Thus, within the ICC framework, a Laplace approximation to the posterior distribution is adopted due to its relative speed and ease to compute.
Laplace’s approximation provides an analytical expression for the posterior distribution. Define $J(\boldsymbol{\theta})\coloneqq\log\pi(\boldsymbol{\theta}\>|\>\boldsymbol{y})$, and let $\hat{\boldsymbol{\theta}}$ be the parameter vector which maximizes $J(\boldsymbol{\theta})$, $\hat{\boldsymbol{\theta}}=\operatorname*{arg\,max}_{\boldsymbol{\theta}\in\boldsymbol{\Theta}}J(\boldsymbol{\theta})$—also known as the maximum a posteriori (MAP) estimate. The covariance of the approximated posterior is determined by the inverse of the Hessian, $\boldsymbol{H}$, of $J(\hat{\boldsymbol{\theta}})$, $\Sigma=\boldsymbol{H}_{J}(\hat{\boldsymbol{\theta}})^{-1}$. In this work, the Hessian is calculated through central finite difference. Thus, the Laplace approximation to the posterior is a Gaussian distribution characterized by the MAP parameter estimate and the local curvature at the MAP estimate, $\tilde{\pi}^{L}(\boldsymbol{\theta}\>|\>\boldsymbol{y})=\mathcal{N}_{D}(\hat{\boldsymbol{\theta}},\Sigma)$. This approach for approximating the posterior is very efficient and makes quasi-real-time step selection and model calibration a viable task.
@@ 3.4.2 Bayesian Optimal Experimental Design
The ICC framework actively drives the experiment through the load path that will yield informative calibration data by leveraging Bayesian optimal experimental design (BOED). Let $\boldsymbol{\xi}$ denote the load path of the specimen from start to finish and $\xi_{t}$ denote each load step from $t=1,\dots,T$, where $T$ is the total number of load steps in the load path ($T=5$ in this example). Also, let $\boldsymbol{y}_{t}$ be the data which is collected at load step $t$.
Within each cycle of the ICC framework, Bayesian inference is performed (Fig. 1d), which produces a posterior probability distribution that quantifies the parameter uncertainty conditioned on the data collected thus far. For example, after collecting data at load step $t$, the posterior is obtained via $\pi(\boldsymbol{\theta}\>|\>\boldsymbol{y}_{1:t},\boldsymbol{\xi}_{1:t})\propto f(\boldsymbol{y}_{1:t}\>|\>\boldsymbol{\theta},\boldsymbol{\xi}_{1:t})\pi(\boldsymbol{\theta})$. Next, load step $\xi_{t+1}$ must be selected (Fig. 1e). While this selection has traditionally been accomplished through subject matter expertise, BOED accomplishes this task in the ICC framework by taking into account previously collected data and the current state of knowledge about the unknown parameters to determine which load step, from a set of candidate load steps $\Xi$, will be the most informative.
Specifically, in this work, the decision is made through the calculation of the expected information gain (EIG), which is derived from the Kullback-Leibler (KL) divergence. The KL divergence provides a measure of the statistical distance between two probability distributions; in this setting, the two distributions of interest are the prior $\pi(\boldsymbol{\theta})$ and the posterior $\pi(\boldsymbol{\theta}\>|\>\boldsymbol{y},\xi)$ after performing candidate load step $\xi$. The candidate load steps are the branches in the load path tree, which here are applying either a displacement increment along Axis A (horizontal) or Axis B (vertical) of the cruciform specimen, $\xi\in\Xi=\{\text{Axis A},\text{Axis B}\}$. The optimal load step is then the one that has the greater expected KL divergence, which means it is expected to yield a posterior distribution that has the greater statistical distance from the prior, and therefore contains more information about the parameters. Since the KL divergence is calculated before the experiment is performed and the data collected, in practice, the expectation of the KL divergence over the marginal distribution of all possible outcomes is used to yield the EIG of candidate load step $\xi$,
@@
The approximation in Eqn. (9) uses samples from the likelihood and prior to evaluate the inner and outer sum of the Monte Carlo estimator. The quality of the estimate is largely determined by $N$ and $M$, the number of samples used to evaluate the outer and inner sums, which control the estimator variance and bias, respectively [63].
@@
In order to incorporate previously collected data into the EIG calculation, at each decision point, the prior in Eqn. (9) is defined as the posterior from the previous step. For example, when choosing load step $t+1$, $\pi(\boldsymbol{\theta})\coloneqq\pi(\boldsymbol{\theta}\>|\>\boldsymbol{y}_{1:t},\xi_{1:t})$, resulting in parameter samples being drawn from the calculated posterior distribution conditioned on data from load steps $1,\dots,t$: $\boldsymbol{\theta}_{i,j}\sim\pi(\boldsymbol{\theta}\>|\>\boldsymbol{y}_{1:t},\boldsymbol{\xi}_{1:t})$. To start the EIG process and determine the first load step from the undeformed state, either (a) the EIG may be calculated for the two candidate load steps by sampling from the prior, or (b) the first load step may be pre-determined, in which case there is no EIG calculation for the first load step. In this work, the second option was pursued. Finally, in this work, a myopic (a.k.a. greedy) BOED algorithm is used, meaning a locally optimal decision is made for each load step, as opposed to a globally optimal decision for the entire load path. This helps preserve computational resources as the EIG can be very costly to compute, and the cost would increase exponentially if calculating for the global optimum.
A challenge of computing the EIG is that numerical underflow may occur as a result of one or more factors: 1) if the prior distribution does not place a sufficient amount of mass on the posterior, leading to calculations for the evidence term (the inner loop of the EIG MC estimator) that are below machine precision, and 2) in the presence of high-dimensional data where the covariance-scaled squared difference of the measurements and model output—the exponential term in Eqn. (13)—is less than 1. The former was encountered and successfully handled through importance sampling (details in C). The latter case is observed if the displacement field is used for calibration without dimension reduction; however, this contributing factor was easily resolved with the PCA dimension reduction process discussed in Sec. 3.2. In combination, these two approaches were successful in resolving issues with underflow.
In this work, $M=1\times 10^{3}$ and $N=1\times 10^{4}$, and the initial load step was chosen to be $\xi_{1}=\text{Axis A}$. Pre-determining the first load step preserved computational resources as this reduced the number of surrogate models that needed to be built: they were only built for the portion of the tree which followed Node 1 in Fig. 2. Additionally, as was observed in [27], if there is no previously collected data (i.e. prior to the first load step), the EIG is determined solely by the prior modeling choice and physical model. So in cases such as this where the prior is only weakly informative, the EIG is unlikely to show strong preference between the two load steps. After the initial load step, the EIG was calculated to determine all subsequent load steps in the load path.
@@ 4 Synthetic Cruciform Exemplar
In this section, the exemplar problem is presented and discussed. The statistical model and other algorithmic settings within the ICC framework are discussed (Sec. 4.1), and results of the load path selection and calibration are presented (Sec. 4.2).
@@ 4.1 Problem Setup
The exemplar problem explored in this work is the calibration of a constitutive model for a cruciform specimen made from an aluminum alloy being deformed in a biaxial load frame. The model form under consideration is the elastoplastic constitutive model described in Sec. 3.1 with unknown parameter vector $\boldsymbol{\theta}:=[\sigma_{y},\,A,\,n,\,a]^{T}$. These are the initial yield stress $\sigma_{y}$ and the isotropic hardening variables $A$ and $n$ from Eqn. (4) and the exponent $a$ from Eqn. (3). This section is used to describe the statistical modeling decisions and other algorithm settings specific to this exemplar problem.
The parameters were constrained to take on values within the bounds specified in Table 2, which are the same as those use for surrogate training. The prior model was chosen to be a truncated normal probability distribution (TN) which enforced these bounds during the calibration,
@@
The distribution means $\mu$ were chosen to yield an expected value within the bounds, and the variances $\delta^{2}$ were chosen to be moderately diffuse. The support of the distribution, defined by $lb$ (the lower bound) and $ub$ (the upper bound), was chosen to be the same as that used for the surrogate training. All distribution parameters are detailed in Table 3.
@@
Table 3: Means $\mu_{\theta}$, variances $\delta^{2}_{\theta}$ and bounds $lb$ and $ub$ for the truncated normal prior model Eqn. (11) used in the exemplar problem. Values for parameters $\sigma_{y}$ and $A$ have units of ksi.
@@
Simulated experimental data was generated by running the expensive FEA under assumed true parameter values $\boldsymbol{\theta}^{true}$ (detailed in Table 1) with added Gaussian error. This synthetic data is used here for model calibration to demonstrate the ICC framework under controlled conditions in preparation for later using it for real, physical experiments. The data consisted of the nodal displacements, dimensionally reduced via PCA to yield the singular values, and global load values for both the $X$ and $Y$ components, $\boldsymbol{y}=[\text{disp}^{\text{PCA}}_{i},\text{load}_{i}],\,i\in[X,Y]$, where $\text{disp}^{\text{PCA}}_{i}=\text{disp}_{i}V^{*}_{i}$ and $V^{*}_{i}$ is the matrix containing the retained principal components for directional component $i$. The FEA model was evaluated and measurements collected at the end of each load step, corresponding to the nodes in the load path tree.
@@
The data generation process was modeled under the assumption that the observed data is equal to the surrogate model output (which consists of the dimensionally reduced nodal displacements and load values) with added Gaussian error $e$,
@@
$\boldsymbol{y}_{h,i}=\tilde{g}(\boldsymbol{\theta})_{h,i}+e_{h,i},\qquad e_{h,i}\sim\mathcal{N}_{N_{h,i}}\left(0,\Psi_{h,i}\right),\,h\in[\text{disp}^{\text{PCA}},\text{load}],\,i\in[X,Y],$ (12)
@@
where $\boldsymbol{y}_{h,i}$ and $\tilde{g}(\boldsymbol{\theta})_{h,i}$ are the data the surrogate replacement of the forward model for each QoI and directional component evaluated at $\boldsymbol{\theta}$, and $N_{h,i}$ is the number of measured observations for each QoI and directional component: $N_{\text{disp}^{\text{PCA}}}=5$ and $N_{\text{load}}=1$—the same for both $X$ and $Y$. The equivalent likelihood function for the expression in Eqn. (12) is a multivariate normal distribution centered at the surrogate replacement of the forward model with a covariance matrix $\Psi_{h,i}$ of size $N_{h,i}\times N_{h,i}$,
@@
$f\left(\boldsymbol{y}_{h,i}\>|\>\boldsymbol{\theta},\boldsymbol{\xi}\right)$ $=\mathcal{N}_{N_{h,i}}\left(\boldsymbol{y}_{h,i};\tilde{g}(\boldsymbol{\theta})_{h,i},\Psi_{h,i}\right),$ (13)
$=(2\pi)^{-N_{h,i}/2}\det\left(\Psi_{h,i}\right)^{-1/2}\exp\left(-\frac{1}{2}\left(\boldsymbol{y}_{h,i}-\tilde{g}(\boldsymbol{\theta})_{h,i}\right)^{T}\Psi_{h,i}\left(\boldsymbol{y}_{h,i}-\tilde{g}(\boldsymbol{\theta})_{h,i}\right)\right),$
$h\in[\text{disp}^{\text{PCA}},\text{load}],\,i\in[X,Y].$
@@
The observation noise variance is assumed to be independent and identically distributed (i.i.d.) such that $\Psi_{h,i}=\psi^{2}_{h,i}\mathbb{I}_{N_{h,i}}$, where $\mathbb{I}_{N_{h,i}}$ is an $N_{h,i}\times N_{h,i}$ identity matrix and $\psi^{2}_{h,i}$ is the observation noise variance—assumed to be known and fixed at $\psi^{2}_{\text{disp}}=4\mu\text{m}^{2}$ and $\psi^{2}_{\text{load}}=582.11\,\mbox{N}^{2}$ for both the $X$ and $Y$ directional components. Note, the experimental noise variance for the displacement was defined in the field space and then transformed to account for the linear transformation of the displacement field via PCA, $\Psi_{\text{disp}^{\text{PCA}},i}=V^{*T}_{i}\Psi_{\text{disp},i}V^{*}_{i},\,i\in[X,Y]$—additional details may be found in D. The noise level was chosen to be on the same order as that expected in experimental data. The likelihood is written as a product over the 4 data sets contained in $\boldsymbol{y}$ (load and displacement for the $X$ and $Y$ directional components) under the assumption of independence,
@@
$f\left(\boldsymbol{y}\>|\>\boldsymbol{\theta},\boldsymbol{\xi}\right)=\prod_{h}\prod_{i}f\left(\boldsymbol{y}_{h,i}\>|\>\boldsymbol{\theta},\boldsymbol{\xi}\right),\quad h\in[\text{disp}^{\text{PCA}},\text{load}],\quad i\in[X,Y].$ (14)
@@
The ICC algorithm was carried out to a total of $T=5$ load steps, with each load step applying a displacement increment of 0.25 mm along the specified axis. The first load step was pre-determined to be $\xi_{1}=\text{Axis A}$. After the initial load step, the EIG was calculated to determine all subsequent load steps in the load path—$[\xi_{2},\xi_{3},\xi_{4},\xi_{5}]$.

### v3

@@ 3.4 Bayesian Techniques
A brief discussion of the Bayesian paradigm for inference and experimental design is contained in this section. The reader is referred to [30] for further discussion on the theory and tools found in this work.
@@ 3.4.1 Bayesian Inference
In each cycle of the ICC feedback loop, a single load step is applied to the specimen, and the collected data is used to perform model calibration. In lieu of a deterministic approach for model calibration, Bayesian inference is used, which provides an elegant way to quantify parameter uncertainty. In this work, inference was performed with both load and displacement field measurements, although it could have been performed using the dimensionally-reduced displacement measurements in the PCA space as well; see Sec. 5.
The first step is to model any knowledge regarding the unknown parameters of the forward model, $\boldsymbol{\theta}\in\boldsymbol{\Theta}\subset\mathbb{R}^{D}$, that is available a priori to observing data. In the adopted notation, the parameters $\boldsymbol{\theta}$ belongs to subspace $\boldsymbol{\Theta}$ of dimension $D$. Data $\boldsymbol{y}\in\boldsymbol{Y}\subset\mathbb{R}^{N_{obs}}$ is then collected which belongs to the data space $\boldsymbol{Y}$ with dimensionality $N_{obs}$, which is the number of measurements. The likelihood function is a probability model that describes the model fitness to the data (or the likelihood of observing the data) for a given parameter value, $f(\boldsymbol{y}\>|\>\boldsymbol{\theta})$. Once this data become available, the prior model is conditioned upon it and updated to the posterior distribution, $\pi(\boldsymbol{\theta}\>|\>\boldsymbol{y})$, via Bayes’ rule,
@@
Thus, both the data and prior beliefs are taken into account when calculating the posterior distribution, which is, generally speaking, not available in closed-form due to the dimensionality of the parameter space. For this reason, simulation techniques such as Markov Chain Monte Carlo (MCMC) [63, 64, 65, 66, 67] are widely used to obtain an approximation of the posterior distribution through one or more of a variety of sampling algorithms. In the ICC framework, the posterior distribution must be approximated after each load step, with the total number of load steps predetermined to be 5 in this case. As the goal is to eventually use this method in real-time decision making with experiments, less than an hour is allotted to posterior estimation within each cycle of the ICC framework in order to complete the experiment (including setup time) in one day. [^2] This time allotment would reduce with additional load step options or a larger number of total load steps. Therefore, in this application of real-time decision-making, the expense of MCMC would be challenging to overcome. Thus, within the ICC framework, a Laplace approximation to the posterior distribution is adopted due to its relative speed and ease to compute.
Laplace’s approximation, based on a Gaussian assumption, provides an analytical expression for the posterior distribution. Define $J(\boldsymbol{\theta})\coloneqq\log\pi(\boldsymbol{\theta}\>|\>\boldsymbol{y})$, and let $\hat{\boldsymbol{\theta}}$ be the parameter vector which maximizes $J(\boldsymbol{\theta})$, $\hat{\boldsymbol{\theta}}=\operatorname*{arg\,max}_{\boldsymbol{\theta}\in\boldsymbol{\Theta}}J(\boldsymbol{\theta})$—also known as the maximum a posteriori (MAP) estimate. The covariance of the approximated posterior is determined by the inverse of the Hessian, $\boldsymbol{H}$, of $J(\hat{\boldsymbol{\theta}})$, $\Sigma=\boldsymbol{H}_{J}(\hat{\boldsymbol{\theta}})^{-1}$. In this work, the Hessian is calculated through central finite difference. Thus, the Laplace approximation to the posterior is a Gaussian distribution characterized by the MAP parameter estimate and the local curvature at the MAP estimate, $\tilde{\pi}^{L}(\boldsymbol{\theta}\>|\>\boldsymbol{y})=\mathcal{N}_{D}(\hat{\boldsymbol{\theta}},\Sigma)$. This approach for approximating the posterior is very efficient and makes quasi-real-time step selection and model calibration a viable task. An MCMC simulation was performed and included in B to confirm that the posterior is well-approximated by a Gaussian distribution in this example.
@@ 3.4.2 Bayesian Optimal Experimental Design
The ICC framework actively drives the experiment through the load path that will yield informative calibration data by leveraging Bayesian optimal experimental design (BOED). Let $\boldsymbol{\xi}$ denote the load path of the specimen from start to finish and $\xi_{t}$ denote each load step from $t=1,\dots,\mathcal{T}$, where $\mathcal{T}$ is the total number of load steps in the load path ($\mathcal{T}=5$ in this example). Also, let $\boldsymbol{y}_{t}$ be the data which is collected at load step $t$.
Within each cycle of the ICC framework, Bayesian inference is performed (Fig. 1(c)), which produces a posterior probability distribution that quantifies the parameter uncertainty conditioned on the data collected thus far. For example, after collecting data at load step $t$, the posterior is obtained via $\pi(\boldsymbol{\theta}\>|\>\boldsymbol{y}_{1:t},\boldsymbol{\xi}_{1:t})\propto f(\boldsymbol{y}_{1:t}\>|\>\boldsymbol{\theta},\boldsymbol{\xi}_{1:t})\pi(\boldsymbol{\theta})$. Next, load step $\xi_{t+1}$ must be selected (Fig. 1(d)). While this selection has traditionally been accomplished through subject matter expertise, BOED accomplishes this task in the ICC framework by taking into account previously collected data and the current state of knowledge about the unknown parameters to determine which load step, from a set of candidate load steps $\Xi$, will be the most informative.
Specifically, in this work, the decision is made through the calculation of the expected information gain (EIG), which is derived from the Kullback-Leibler (KL) divergence. The KL divergence provides a measure of the statistical distance between two probability distributions; in this setting, the two distributions of interest are the prior $\pi(\boldsymbol{\theta})$, which contains all the information up to this point (details following), and the posterior $\pi(\boldsymbol{\theta}\>|\>\boldsymbol{y},\xi)$ after performing candidate load step $\xi$. The candidate load steps are the branches in the load path tree, which here are applying either a displacement increment along the X (horizontal) axis (load step A) or the Y (vertical) axis (load step B) of the cruciform specimen, $\xi\in\Xi=\{\text{A},\text{B}\}$. The preferred load step is then the one that has the greater expected KL divergence, which means it is expected to yield a posterior distribution that has the greater statistical distance from the prior and therefore contains more information about the parameters. Since the KL divergence is calculated before the experiment is performed and the data collected, in practice, the expectation of the KL divergence over the marginal distribution of all possible outcomes is used to yield the EIG of candidate load step $\xi$,
@@
The approximation in Eqn. (9) uses samples from the likelihood and prior to evaluate the inner and outer sum of the Monte Carlo estimator. The quality of the estimate is largely determined by $N$ and $M$, the number of samples used to evaluate the outer and inner sums, which control the estimator variance and bias, respectively [68]. In order avoid numerical underflow, the log-sum-exp of the log likelihoods is used to stabilize the inner sum [69].
@@
In order to incorporate previously collected data into the EIG calculation, at each decision point, the prior in Eqn. (9) is defined as the posterior from the previous step. For example, when choosing load step $t+1$, $\pi(\boldsymbol{\theta})\coloneqq\pi(\boldsymbol{\theta}\>|\>\boldsymbol{y}_{1:t},\xi_{1:t})$, resulting in parameter samples being drawn from the calculated posterior distribution conditioned on data from load steps $1,\dots,t$: $\boldsymbol{\theta}_{i,j}\sim\pi(\boldsymbol{\theta}\>|\>\boldsymbol{y}_{1:t},\boldsymbol{\xi}_{1:t})$. To start the EIG process and determine the first load step from the undeformed state, either (a) the EIG may be calculated for the two candidate load steps by sampling from the prior, or (b) the first load step may be pre-determined, in which case there is no EIG calculation for the first load step. In this work, the second option was pursued. Finally, in this work, a myopic (a.k.a. greedy) BOED algorithm is used, meaning a local or short-sighted decision is made for each load step, as opposed to a global decision which considers the entire load path. This helps preserve computational resources as the EIG can be very costly to compute, and the cost would increase exponentially if calculating for the globally preferred path. Searching for the globally preferred full load path may be made possible with the use of a sequential OED approach, as was done in [70]; however, exploration of this approach is reserved for future endeavors as sequential OED would be particularly challenging for this application due to the non-Markovian nature of materials.
In this work, $M=1\times 10^{3}$ and $N=1\times 10^{4}$, and the initial load step was chosen to be $\xi_{1}=\text{A}$. Pre-determining the first load step preserved computational resources as this reduced the number of surrogate models that needed to be built: they were only built for the portion of the tree which followed Node 1 in Fig. 2. Additionally, as was observed in [30], if there is no previously collected data (i.e. prior to the first load step), the EIG is determined solely by the prior modeling choice and physical model. So in cases such as this where the prior is only weakly informative, the EIG is unlikely to show strong preference between the two load steps. After the initial load step, the EIG was calculated to determine all subsequent load steps in the load path.
@@ 4 Synthetic Cruciform Exemplar
In this section, the exemplar problem is presented and discussed. The statistical model and other algorithmic settings within the ICC framework are discussed (Sec. 4.1) and results of the load path selection and calibration are presented (Sec. 4.2).
@@ 4.1 Problem Setup
The exemplar problem explored in this work is the calibration of a constitutive model for a cruciform specimen made from an aluminum alloy being deformed in a biaxial load frame. The model form under consideration is the elastoplastic constitutive model described in Sec. 3.1 with unknown parameter vector $\boldsymbol{\theta}:=[\sigma_{y},\,A,\,n,\,a]^{T}$. These are the initial yield stress $\sigma_{y}$ and the isotropic hardening variables $A$ and $n$ from Eqn. (4) and the exponent $a$ from Eqn. (3).
The parameters were constrained to take on values within the bounds specified in Table 2, which are the same as those use for surrogate training. The prior model for each parameter was chosen to be a truncated normal probability distribution (TN) which enforced these bounds during the calibration,
@@
The distribution means $\mu$ were chosen to yield an expected value within the bounds, and the variances $\delta^{2}$ were chosen to be moderately diffuse. The support of each distribution, defined by $lb$ (the lower bound) and $ub$ (the upper bound), was chosen to be the same as that used for the surrogate training. All distribution parameters are detailed in Table 3 for each parameter. The truncated normal prior distributions were only mildly informative, showing just a slight preference for the mean of the bounding region; a comparative analysis using a uniform prior yielded results with no noticeable difference.
@@
Table 3: Means $\mu_{\theta}$, variances $\delta^{2}_{\theta}$ and bounds $lb$ and $ub$ for the truncated normal prior model on each parameter, Eqn. (11), used in the exemplar problem. Values for parameters $\sigma_{y}$ and $A$ have units of ksi.
@@
Simulated experimental data was generated by running the expensive FEA under assumed true parameter values $\boldsymbol{\theta}^{true}$ (detailed in Table 1) with added i.i.d. Gaussian error. This synthetic data is used here for model calibration to demonstrate the ICC framework under controlled conditions in preparation for later using it for real, physical experiments. The data consisted of the nodal displacements and global load values for both the $X$ and $Y$ components, $\boldsymbol{y}=[\text{disp}_{i},\text{load}_{i}],\,i\in[X,Y]$. The FEA model was evaluated and measurements collected at the end of each load step, corresponding to the nodes in the load path tree.
@@
The data generation process was modeled under the assumption that the observed data is equal to the surrogate model output with added Gaussian error $e$,
@@
$\boldsymbol{y}_{hi}=\tilde{g}(\boldsymbol{\theta})^{r}_{hi}+e_{hi},\qquad e_{hi}\sim\mathcal{N}_{N_{hi}}\left(0,\Psi_{hi}\right),\,h\in[\text{disp},\text{load}],\,i\in[X,Y],$ (12)
@@
where $\boldsymbol{y}_{hi}$ and $\tilde{g}(\boldsymbol{\theta})^{r}_{hi}$ are the data the surrogate replacement of the forward model for each QoI and directional component evaluated at $\boldsymbol{\theta}$. The $r$ superscript for the surrogate output indicates that the full-field displacements were reconstructed from the predicted surrogate model PCA singular values via the process described in Sec. 3.2. $N_{hi}$ is the number of measured observations for each QoI and directional component: $N_{\text{disp}}=1,022$ and $N_{\text{load}}=1$—the same for both $X$ and $Y$. The equivalent likelihood function for the expression in Eqn. (12) is a multivariate normal distribution centered at the surrogate replacement of the forward model with a covariance matrix $\Psi_{hi}$ of size $N_{hi}\times N_{hi}$,
@@
$f\left(\boldsymbol{y}_{hi}\>|\>\boldsymbol{\theta},\boldsymbol{\xi}\right)=\mathcal{N}_{N_{hi}}\left(\boldsymbol{y}_{hi};\tilde{g}(\boldsymbol{\theta})^{r}_{hi},\Psi_{hi}\right)\qquad h\in[\text{disp},\text{load}],\,i\in[X,Y].$ (13)
@@
The observation noise variance is assumed to be independent and identically distributed (i.i.d.) such that $\Psi_{hi}=\psi^{2}_{hi}\mathbb{I}_{N_{hi}}$, where $\mathbb{I}_{N_{hi}}$ is an $N_{hi}\times N_{hi}$ identity matrix and $\psi^{2}_{hi}$ is the observation noise variance—assumed to be known and fixed at $\psi^{2}_{\text{disp}}=4$µm2 and $\psi^{2}_{\text{load}}=582.11\,\mbox{N}^{2}$ for both the $X$ and $Y$ directional components. The value chosen for the load measurement noise equates to a standard deviation of approximately 0.05% of the highest expected force at the final load step predicted from simulations with a nominal set of parameters. Likewise, the value chosen for the displacement measurement noise mimics the expected noise of DIC displacement measurements, based on an assumed DIC noise of 0.05 px (typical of a well-controlled DIC measurement in a laboratory environment), a field-of-view of 100 mm and a 5 MP camera with resolution of $2048\times 2448$ px2. The likelihood is written as a product over the 4 data sets contained in $\boldsymbol{y}$ (load and displacement for the $X$ and $Y$ directional components) under the assumption of independence,
@@
$f\left(\boldsymbol{y}\>|\>\boldsymbol{\theta},\boldsymbol{\xi}\right)=\prod_{h}\prod_{i}f\left(\boldsymbol{y}_{hi}\>|\>\boldsymbol{\theta},\boldsymbol{\xi}\right),\quad h\in[\text{disp},\text{load}],\quad i\in[X,Y].$ (14)
@@
The ICC algorithm was carried out to a total of $\mathcal{T}=5$ load steps, with each load step applying a displacement increment of 0.25 mm along the specified axis. The first load step was pre-determined to be $\xi_{1}=\text{A}$. After the initial load step, the EIG was calculated to determine all subsequent load steps in the load path—$[\xi_{2},\xi_{3},\xi_{4},\xi_{5}]$.